OpenStudy (anonymous):

3 2/5 * 2 2/3 plz help!!!!

OpenStudy (anonymous):

\[ 3\dfrac{2}{5}\cdot2\dfrac{2}{3}\\ \dfrac{17}{5}\cdot\dfrac{8}{3}\\ \dfrac{17\cdot8}{5\cdot3}\\ \dfrac{136}{15} \]